LSL R3, R3, R12

Memory

Address

Orig.

Data

Updated

Data

Register

Orig.

Data

Updated

Data

0x84F0

0x82

R0

0xDC002B45

0x84F1

0x32

R1

0x00008500

0x84F2

0xBC

R2

0xAB03873C

0x84F3

0xA4

R3

0x2B78DFE0

0x84F4

0xF4

R4

0x000084F0

0x84F5

0x6A

R5

0x00100000

0x84F6

0x9D

R6

0x041B0C0D

0x84F7

0xC2

R7

0x00100000

0x84F8

0x42

R8

0x00000000

0x84F9

0x67

R9

0x00842AB0

0x84FA

0x8D

R10

0x0ABD9830

0x84FB

0xA1

R11

0x0000000A

0x84FC

0xF2

R12

0x00000002

0x84FD

0xB8

R13

0X00008500

0x84FE

0x5E

R14

0x003762DD

0x84FF

0xEE

R15

0x00100000

0x8500

0xFA

0x8501

0xF3

Flags

0x8502

0x54

N

0

0x8503

0xBA

Z

1

0x8504

0x6D

C

1

0x8505

0x9C

V

0

0x8506

0x78

Q

0

0x8507

0xBB

0x8508

0xC3

0x8509

0xB4

0x850A

0xAA

0x850B

0xD9

0x850C

0xDA

0x850D

0xFF

0x850E

0xE5

0x850F

0x33
Answer #1

The LSL Command left shifts the input in register R3 by a shift amount specified by R12(2)

Shifting by k bits means a multiplication of 2^k.

Value in R3 = 2B78DFE0 in hex = 729341920 in decimal

729341920 * (2^2) = 729341920 * 4 = 2917367680 = 0xADE37F80

Hence the only updated value is in register R3 = 0xADE37F80

No change in any other value.

P.S. Please give a thumbs up if my answer could be of help to you!